OpenStudy (anonymous):

What is the probability of rolling a 6 and then a 1 on consecutive rolls of a standard, 6-sided number cube?

OpenStudy (anonymous):

the probability of each is \(\frac{1}{6}\) so the probability of both occurring is \[\left(\frac{1}{6}\right)^2\]